What is the maximum possible concentration of Ba2+ ions in a NaOH solution if the concentration of OH is maintained at 0.100 M? The solubility product constant of Ba(OH)2 is Ksp = 5.0 × 10–3.

  •  A. 0.32 M
  •  B. 0.020 M
  •  C. 0.50 M
  •  D. 0.080 M
  •  E. 0.125 M
1
Expert's answer
2021-03-10T06:15:04-0500

Solution:

NaOH → Na+ + OH

According to the equation above: C(NaOH) = [Na+] = [OH]

Hence, [OH] = C(NaOH) = 0.100 M


Ba(OH)2 → Ba2+ + 2OH

Ba(OH)2 contains 2 hydroxide ions (OH), therefore: [OH] = 2 × 0.100 M = 0.200 M

Ksp = [Ba2+] × [OH]2

5.0×10–3 = [Ba2+] × (0.200)2

[Ba2+]max = (5.0×10–3) / (0.04) = 0.125 M

[Ba2+]max = 0.125 M


Answer: E. 0.125 M